# Day-Ahead Electricity Price Forecasting Using Merit-Order Curves Time Series

### v1

@@ 4.1 Curves definition
From the definition we can notice that merit-order curves are non-continu-ous monotone (increasing for supply and decreasing for demand) step functions. A direct consequence is that admits an inverse which is itself a monotone step function. It is interesting to note that this inverse has an easier mathematical notation and interpretation. Indeed, its expression matches (for supply curves) the unnormalized666Given the ’s don’t sum to 1. cumulative distribution function (cdf) of a discrete distribution over the support . For demand, the only difference is that it is the reverse cdf (or survival function). Using the superscript for supply and for demand, we have:
That is, is the quantity that is offered (or demanded) at a price below (resp. above) . Note that the definition above does not require the sequence of offers (or bids) to be price-ordered. It is more intuitive to work with the quantity function as it represents a more familiar mathematical object than the price function . The quantity function simply tells us about the total offered quantity and how this offered quantity is distributed on the price domain. It also conveniently changes the perspective on the curves variability which is mainly attributed to their inelastic component: seasonal power demand variations for demand curves and renewable energy sources (RES) generation for supply curves.
@@ 4.2 Curves representation
The goal is to model and forecast the time series of curves and over a time period composed of hourly intervals.
Functional Data Analysis (FDA) is a framework that extends traditional statistical methods to handle data that vary over a continuum, such as curves, surfaces, or other functional forms. Unlike classical approaches that analyze discrete points, FDA treats entire functions as the fundamental unit of analysis. It assumes however that only discrete and noisy samples of continuous functions are observed and therefore the first step of any FDA is to estimate the underlying functions from the data – a process known as smoothing.
In our case, things are a bit different: we observe the full functions and these functions are not continuous. Nevertheless, as we saw that the quantity function is expressed as the (unnormalized) cdf of some discrete distribution, with a non-fixed, random support, it may be interpreted as a realization of a Dirichlet process (ferguson_bayesian_1973; sammut_dirichlet_2011) with a continuous base distribution. The corresponding base distribution has a smooth cdf, which then becomes the natural object of analysis.
@@
In order to forecast functional time series, we follow the guidance of aue_prediction_2015 that recommend the use of functional principal component analysis (FPCA) to derive an uncorrelated vector representation of the curves and apply standard multivariate time series forecasting methods to this representation, for finally obtaining a curves forecast using the Karhunen-Loève (KL) expansion (horvath_inference_2012). FPCA is a dimensionality reduction technique for functional data, extending classical PCA to infinite-dimensional spaces. It decomposes a set of observed functions into an orthonormal basis that captures the dominant modes of variations in the data. This decomposition yields the truncated KL expansion:
@@
where is the mean function, are the first functional principal components (FPCs) and are the scores of observation on the first FPCs. Intuitively, the FPCs can be seen as specific features common to all curves while the scores are measures of how pronounced are these features in a specific curve. The perfect equality holds when but in practice a satisfying approximation can be found with a finite , typically chosen – as in traditional PCA – as the elbow point of the scree plot or such that the ratio of explained variance surpasses a certain threshold (johnson_applied_2007).
@@
Since FDA assumes smooth functions, the estimated FPCs should also be smooth functions. This smoothness can be achieved either by incorporating a roughness penalty directly into the FPCA estimation procedure or by pre-smoothing the functional data curves prior to applying FPCA. We adopt the latter approach for computational simplicity, utilizing kernel smoothing with the Nadaraya-Watson kernel estimator (ferraty_nonparametric_2004). The bandwidth parameter of the kernel function, which controls the level of smoothness, is chosen using the generalized cross-validation criterion.
Aue et al. show that a vector autoregressive model of order – VAR() – applied to the -dimensional vector representation given by the scores on the first principal components is asymptotically equivalent to a functional autoregressive model (bosq_linear_2000) of the same order – FAR() – when . Therefore, they argue that forecasting the functional time series by forecasting the vector time series of the FPC scores (given that it provides a satisfying dimensionality reduction) with any multivariate forecasting method – being a VAR model or not – is theoretically justified and is easier than devising complex functional models. Note that the uncorrelation property of the FPC scores has motivated approaches where univariate models are separately fitted to each of the score series (hyndman_robust_2007; hyndman_forecasting_2009). As highlighted by aue_prediction_2015, though it can turn out to be a simple and efficient way to forecast the vector of scores, it has the major drawback of overlooking the cross-correlation that can exist at lags greater than zero, since the FPCA procedure only guarantees that the vectors have no instantaneous (lag 0) correlation.
@@
The advantage of working with is triple: (i) we turn an infinite-dimension problem into an intrinsically finite-dimension one, (ii) we phase out the intrinsic dependence of functional data, (iii) we jointly model the supply and demand processes.
We forecast with some multivariate forecasting model (detailed in the next section), with possibly a set of exogenous covariates . We then inverse-transform a -step-ahead forecast using the truncated KL expansion to get a forecast for the curves pair .
It can happen that curves predictions are not perfectly monotonic. Though preliminary experiments suggested the occurence of little deviations only and low impact on clearing price predictions, it is still desirable to respect this structural constraint. To do so, we post-process the curves predictions using isotonic regression with the pool-adjacent violators algorithm (leeuw_isotone_2009). Being a highly efficient algorithm, it has a negligible impact on computation times.
@@ 4.3 Forecasting the curves’ vector representation
We consider here the day-ahead forecasting problem, regarding the vector hourly time series as 24 separate daily time series, one for each hour of the day. This is common practice in EPF as the 24 hours of the next day are simultaneously settled the day before. In addition, market dynamics vary a lot depending on the hour of the day, justifying to treat them separately. For a detailed motivation of this choice, the reader can refer to ziel_day-ahead_2018. We hence change the time indexing of to consider the value at day and hour , and we solve 24 one-step-ahead forecasting problems.
@@
We consider four variants of the popular parameter-rich ARX model estimated with LASSO widely used in the context of electricity price forecasting (ziel_forecasting_2016; uniejewski_automated_2016; ziel_day-ahead_2018; lago_forecasting_2021; uniejewski_regularization_2024). First this framework was shown to be highly performing for price forecasting and second, it was used in ziel_electricity_2016. Specifically, we consider two univariate models treating each component of separately as suggested by hyndman_forecasting_2009 – the concurrent ARX and the full ARX – and two multivariate models jointly modeling the vector , the concurrent VARX and the full VARX. The term ”concurrent” means the target variable at hour is influenced only by past values from the same hour . Conversely, ”full” means the variable is also influenced by past values observed at different hours. The comparison between the univariate and the multivariate approaches allow to test the added value of modeling the cross-dependence between scores while the comparison between the concurrent and full approaches allow to test the added value of modeling the cross-dependence between hours.
@@
The *AR* part refers to the fact that lagged values are used to predict future values. In our case, similarly to state-of-the-art EPF models (lago_forecasting_2021) we include lags 1, 2, 3 and 7, considering therefore information up to 7 days before. Additionally, each model has the suffix *X which means that it makes use of exogenous variables which refer to day and available on (i.e., they are known in anticipation or they are themselves one-day-ahead forecasts). This set of exogenous variables is represented by the -dimensional vector . Again, like lago_forecasting_2021, we also include lags 1 and 7 of these exogenous variables, and a three-dimensional vector of dummy variables flagging the day type: Mondays, Working days (from Tuesday to Friday), Saturdays and Holidays (Sundays and bank holidays)777Note this is slightly different from lago_forecasting_2021 who consider one dummy for each day of the week but similar to mestre_forecasting_2020 and addresses the calendar effects issues raised by ziel_electricity_2016..
@@
Finally, still following Lago et al.’s guidance, the -regularization parameter is selected as that minimizing the Akaike information criterion (AIC), exploiting the least angle regression (LARS) algorithm (efron_least_2004) for the parameter search. Once is selected, the model is estimated using the traditional coordinate descent algorithm (tibshirani_regression_1996).
4.3.1 Concurrent ARX (ARX)
The concurrent ARX888Note that this model is equivalent to a , with the difference that it is estimated with the LASSO and not with maximum likelihood. models each component of as:
@@ 4.3.1 Concurrent ARX (ARX)
4.3.2 Full ARX (fARX)
The full ARX999Similarly, this model is equivalent to a VARX(7), applied to the 24-dimensional vector – with lags from 4, 5, 6 constrained to be null – with the difference that it is estimated with the LASSO and not with maximum likelihood., very similar to its homonym in uniejewski_automated_2016, models each component of as:
@@ 4.3.2 Full ARX (fARX)
4.3.3 Concurrent VARX (VARX)
@@
4.3.4 Full VARX (fVARX)
@@
Where we impose the restriction that the off-diagonal terms of must be zero whenever , that is, cross-dependence is allowed between hours, but only within the same component101010In that sense, ”semi-full VARX” could be a more precise denomination but we keep ”full VARX” for simplicity.. This restriction is imposed to have a manageable number of parameters. Indeed, without this condition we would have parameters for each component of while with this condition the number of parameters is only . For instance, for and , the condition implies 147 parameters instead of 975.
@@ 4.4 Clearing price point and probabilistic forecasting
The clearing price point prediction is simply obtained by determining the zero111111Computationally speaking, we take the linear interpolation of the two prices concerned by the sign change. of the function .
@@
The choice of the calibration window is a classic trade-off as short windows have the advantage of capturing potential non-stationarity and correcting eventual forecasts bias while long windows will allow for a more accurate estimation of the errors distribution. To get ”the best of both worlds”, we proceed as in lipiecki_postprocessing_2024: We consider four different window sizes: 28 days (4 weeks), 56 days (8 weeks), 91 days (13 weeks) and 182 days (26 weeks) yielding four cdf estimates . We then compute an ensemble prediction considering vertical (probability) averaging (lichtendahl_is_2013):
Regarding step (1), we estimate the multivariate distribution of the -dimensional forecast error vector . To robustly estimate this distribution which we assume stationary within the calibration window, we consider hour-dependent variances for the marginals but a shared correlation structure: We assume the errors for a given hour are independent and identically distributed (i.i.d.) according to some -dimensional distribution with mean and a covariance matrix which can be decomposed as follows:
@@ 4.5.1 Naive model
4.5.2 ziel_electricity_2016
While our approach differs from that of ziel_electricity_2016 regarding the curves vector representation, the multivariate forecasting techniques applied to that representation are essentially the same. For that reason, to guarantee a fair comparison between the two approaches, we apply the models detailed in the previous section to (i) our vector of supply and demand FPCA scores (ii) the vector of supply and demand price class volumes obtained with the method described in ziel_electricity_2016. We invite the reader to refer to the authors’ paper for a rigorous and detailed description of this procedure (as well as for curves reconstruction), which we refer to as Ziel-Steinert Transformation (ZST), as opposed to our functional principal component analysis (FPCA).
@@
4.5.3 Price-based point forecasting models
We consider three benchmark price-based clearing price forecasting models: ARX, fARX and the Lasso-Estimated AutoRegressive model (LEAR) of lago_forecasting_2021. The first two are already described in the previous section (equations 1 and 2). When both curve-based and price-based models are involved, to distinguish the clearing price forecasting model deriving from the curves model (curve-based) from that directly forecasting the price (price-based), we prefix the curve-based models with the curves representation used, e.g, FPCA-ARX or ZST-ARX.
Finally the LEAR model is very similar to fARX with the difference that it adds cross-hour dependence in the exogenous variables effects and accepts only two of them (usually the load and combined wind-solar generation day-ahead forecasts):
As in lago_forecasting_2021, for all three models, prices are transformed with median-based scaling and the area hyperbolic sine variance stabilizing transformation.
4.5.4 Price-based probabilistic forecasting models
We test the three (price-based) clearing price probabilistic forecasting methods of lipiecki_postprocessing_2024, using the Julia package PostForecasts.jl (lipiecki_postforecastsjl_2025) developed by the authors: Quantile Regression Machine (QRM), Conformal Prediction (CP), Isotonic Distributional Regression (IDR). We also include the naive Normal (N) benchmark121212Note that differently from lipiecki_postprocessing_2024 paper, we consider the version which does not suppose zero-mean errors and therefore accounts for potential forecasts bias.. As these methods are based on post-processing of point predictions, we use the best performing price-based model as a point predictor. For more information, we invite the reader to directly refer to lipiecki_postprocessing_2024.
@@ 4.6 Forecasts evaluation
4.6.1 Curves forecasts
As suggested in ramsay_functional_2005, we measure curves forecasting accuracy with the squared correlation function:
@@ 4.6.1 Curves forecasts
[…]
4.6.2 Clearing price forecasts
We evaluate point predictions with the mean absolute error (MAE), the root mean squared error (RMSE) and the relative MAE (rMAE) as defined in lago_forecasting_2021.
@@ 4.6.2 Clearing price forecasts
Probabilistic forecasting performance is composed of a reliability (or calibration) and sharpness (or resolution141414The two concepts are equivalent when probabilistic forecasts have perfect reliability.) component (nowotarski_recent_2018). As any proper scoring rule, the CRPS presents the advantage of simultaneously evaluating both but it can be interesting to understand how models perform on each of these two components. We assess therefore reliability with the probability integral transform (gneiting_probabilistic_2014)
Similarly to the curves forecasts, we perform DM tests for both point and probabilistic forecasts considering average and per-hour daily errors. For point forecasting we consider the absolute () errors as in lago_forecasting_2021 while we use the CRPS, as was done in nowotarski_recent_2018, for probabilistic forecasting.
@@ 4.7 Daily recalibration
As advised by lago_forecasting_2021 and in order to simulate a live forecasting setting, the entire modeling pipeline is recalibrated on a daily basis using a rolling window of 364 days (52 weeks), since year appears to the best single-window choice for EPF according to marcjasz_selection_2018 and hubicka_note_2019. This means that for every day of the testing period:
@@
For FPCA, we combine the elbow and threshold methods (see section 4.2) by taking the maximum of the two suggested number of FPCs to retain, considering a ratio of 99% for the threshold method. While automating the latter is straightforward, the elbow method relies on a qualitative visual assessment of the scree plot and requires therefore a more sophisticated approach. To automatically find this elbow point in the recalibration procedure, we use the popular knee-point detection method of satopaa_finding_2011.
For ZST, ziel_electricity_2016 do not provide discussion on their choice for the number of price classes to consider. Since our goal is to compare the efficiency of the two curves representation strategies, we should ideally match it on a daily basis with the dimension picked by FPCA. For simplicity though, we take fixed and equal to their maximum value observed across the FPCA recalibration procedure, leaving therefore a slight advantage to ZST.
@@
5 Data
5.1 Supply and demand curves
We analyze the supply and demand merit-order curves of the day-ahead electricity market (MGP) of the Italian Power Exchange (IPEX) over the 2023-2024 period. The testing period covers the whole year 2024 (366 days). Note that probabilistic price predictions require at least 182 days of out-of-sample point predictions so the testing period for probabilistic forecasting starts on July 1st, 2024, covering a period of 6 months. The quantity curves are built according to the procedure described in section 4.1, using the DomandaOfferta dataset retrieved from the FTP server of Gestore del Mercato Elettrico (GME) – the operator of IPEX – which contains all supply offers and demand bids submitted on the MGP. Three points must be highlighted:
5.1.1 Relaxation of transmission constraints
Around 35% of the time, the country-level economic optimum is not feasible due to transmission capacity constraints, which means that the market cannot be cleared from the country-level supply and demand curves. In this case, GME runs an optimization procedure, known as locational marginal pricing (LMP), which splits the entire country-level market pool in a minimum number of ”subpools” of market zones such that, in each subpool, the economic optimum is feasible. As a result, there can be several pairs of supply and demand curves – hence several market clearing prices – for a single hour, one for each subpool. To keep the analysis simple, we did as if the country-level economic optimum was always feasible and built the country-level supply and demand curves for any hour, ignoring the problem of transmission congestion. As a result, the constructed curves reflect the true market clearing 65% of the time, the remaining 35% corresponding to a “virtual” market clearing only. Note the same curves construction procedure was considered in mestre_forecasting_2020. The associated market clearing price is what GME calls the national price without constraints, and though not always being the exact market price for the different Italian bidding zones, is still a key monitored market indicator.
5.1.2 Market coupling
5.1.3 Restriction of the curves domain
Three example pairs of supply and demand curves of the dataset are showed on Fig. 3
5.2 Exogenous predictors
We collect a total of 4 exogenous variables for both curves and price forecasts which are similar to those of mestre_forecasting_2020 and commonly chosen in EPF:
1.
Day-ahead load forecast (1 variable): Country-level day-ahead load forecast obtained from the ENTSO-E Transparency Platform (entso-e_entso-e_2025).
2.
3.
Day-ahead forecast of wind and solar generation (1 variable): Combined country-level day-ahead forecasts of wind and solar generation obtained from a private data provider (lseg_data__analytics_ecmwfgfs_2025), derived from the European Centre for Medium-Range Weather Forecasts (ECMWF) and Global Forecast System (GFS) weather models.

### v2

@@ 2.1 Curves definition
From the definition we can notice that merit-order curves are non-continuous monotone (increasing for supply and decreasing for demand) step functions. A direct consequence is that admits an inverse which is itself a monotone step function. It is interesting to note that this inverse has an easier mathematical notation and interpretation. Indeed, its expression matches (for supply curves) the unnormalized444Given the ’s don’t sum to 1. cumulative distribution function (cdf) of a discrete distribution over the support . For demand, the only difference is that it is the reverse cdf (or survival function). Using the superscript for supply and for demand, we have:
That is, is the quantity that is offered (or demanded) at a price below (resp. above) . Note that the definition above does not require the sequence of offers (or bids) to be price-ordered. It is more intuitive to work with the quantity function as it represents a more familiar mathematical object than the price function . The quantity function simply tells us about the total offered quantity and how this offered quantity is distributed on the price domain. It also conveniently changes the perspective on the curves variability which is mainly attributed to their inelastic component: seasonal power demand variations for demand curves and variable renewable energy generation for supply curves.
@@ 2.2 Curves representation
The goal is to model and forecast the time series of curves and over a time period composed of hourly intervals. Functional data analysis (Ramsay and Silverman, 2005; Ferraty and Vieu, 2006) is a framework that extends traditional statistical methods to handle data that vary over a continuum, such as curves, surfaces, or other functional forms. In order to forecast these functional time series, we follow the guidance of Aue et al. (2015) that recommend the use of functional principal component analysis (FPCA) to derive an uncorrelated vector representation of the curves and apply standard multivariate time series forecasting methods to this representation, finally obtaining a curves forecast using the Karhunen-Loève (KL) expansion (Horváth and Kokoszka, 2012, pp. 37–43). FPCA is a dimensionality reduction technique for functional data, extending classical PCA to infinite-dimensional spaces. It decomposes a set of observed functions into an orthonormal basis that captures the dominant modes of variations in the data. This decomposition yields the truncated KL expansion:
@@
where is the mean function, are the first functional principal components (FPCs) and are the scores of observation on the first FPCs. Intuitively, the FPCs can be seen as specific features common to all curves while the scores are measures of how pronounced are these features in a specific curve. The perfect equality holds when but in practice a satisfying approximation can be found with a finite , typically chosen – as in traditional PCA – as the elbow point of the scree plot or such that the ratio of explained variance surpasses a certain threshold (Johnson and Wichern, 2007, pp. 444–447).
@@
Since FDA assumes smooth functions, the estimated FPCs should also be smooth functions. This smoothness can be achieved either by incorporating a roughness penalty directly into the FPCA estimation procedure or by pre-smoothing the functional data curves prior to applying FPCA. We adopt the latter approach for computational simplicity, utilizing kernel smoothing with the Nadaraya-Watson kernel estimator (Wasserman, 2006, p. 71). The bandwidth parameter of the kernel function, which controls the level of smoothness, is chosen using the generalized cross-validation criterion (Ramsay and Silverman, 2005, p. 97).
@@
We forecast with some multivariate forecasting model (detailed in the next section), with possibly a set of exogenous covariates . We then inverse-transform any -step-ahead forecast using the truncated KL expansion to get a forecast for the curves pair .
It can happen that curves predictions are not perfectly monotonic. Though preliminary experiments suggested the occurence of little deviations only and low impact on clearing price predictions, it is still desirable to respect this structural constraint. To do so, we post-process the curves predictions using isotonic regression with the pool-adjacent violators algorithm (Leeuw et al., 2009). Being a highly efficient algorithm, it has a negligible impact on computation times.
@@ 2.3 Forecasting the curves’ vector representation
We consider here the day-ahead forecasting problem, regarding the vector hourly time series as 24 separate daily time series, one for each hour of the day. This is common practice in EPF as the 24 hours of the next day are simultaneously settled the day before. In addition, market dynamics vary a lot depending on the hour of the day, justifying to treat them separately. For a detailed motivation of this choice, the reader can refer to Ziel and Weron (2018). We hence change the time indexing of to consider the value at day and hour , and we solve 24 one-step-ahead forecasting problems.
@@
We consider four variants of the popular parameter-rich ARX model estimated with LASSO widely used in the context of electricity price forecasting (Ziel, 2016; Uniejewski et al., 2016; Ziel and Weron, 2018; Lago et al., 2021; Uniejewski, 2024). First this framework was shown to be highly performing for price forecasting and second, it was used in Ziel and Steinert (2016). Specifically, we consider two univariate models treating each component of separately as suggested by Hyndman and Shang (2009) – the concurrent ARX and the full ARX (equations 1 and 2) – and two multivariate models jointly modeling the vector , the concurrent VARX and the full VARX (equations 3 and 4). The term ”concurrent” means the target variable at hour is influenced only by past values from the same hour . Conversely, ”full” means the variable is also influenced by past values observed at different hours. The comparison between the univariate and the multivariate approaches allow to test the added value of modeling the cross-dependence between scores while the comparison between the concurrent and full approaches allow to test the added value of modeling the cross-dependence between hours.
@@
The *AR* part refers to the fact that lagged values are used to predict future values. In our case, similarly to state-of-the-art EPF models (Lago et al., 2021) we include lags 1, 2, 3 and 7, considering therefore information up to 7 days before. Additionally, each model has the suffix *X which means that it makes use of exogenous variables which refer to day and available on (i.e., they are known in anticipation or they are themselves one-day-ahead forecasts). This set of exogenous variables is represented by the -dimensional vector . Again, like Lago et al. (2021), we also include lags 1 and 7 of these exogenous variables, and a three-dimensional vector of dummy variables flagging the day type: Mondays, Working days (from Tuesday to Friday), Saturdays and Holidays (Sundays and bank holidays)555Note this is slightly different from Lago et al. (2021) who consider one dummy for each day of the week but similar to Mestre et al. (2020) and addresses the calendar effects issues raised by Ziel and Steinert (2016)..
@@
Finally, still following Lago et al.’s guidance, the -regularization parameter is selected as that minimizing the Akaike information criterion (AIC), exploiting the least angle regression (LARS) algorithm (Efron et al., 2004) for the parameter search. Once is selected, the model is estimated using the traditional coordinate descent algorithm (Tibshirani, 1996).
Concurrent ARX (ARX)
The concurrent ARX models each component of as:
@@
Full ARX (fARX)
The full ARX , very similar to its homonym in Uniejewski et al. (2016), models each component of as:
@@
Concurrent VARX (VARX)
@@
Full VARX (fVARX)
@@
Where we impose the restriction that the off-diagonal terms of must be zero whenever , that is, cross-dependence is allowed between hours, but only within the same component666In that sense, ”semi-full VARX” could be a more precise denomination but we keep ”full VARX” for simplicity.. This restriction is imposed to have a manageable number of parameters. Indeed, without this condition we would have parameters for each component of while with this condition the number of parameters is only . For instance, for and , the condition implies 147 parameters instead of 975.
@@ 2.4 Clearing price point and probabilistic forecasting
The clearing price point prediction is simply obtained by determining the zero777Computationally speaking, we take the linear interpolation of the two prices concerned by the sign change. of the function .
@@
Details for step (1) can be found in A. The choice of the calibration window is a classic trade-off as short windows have the advantage of capturing potential non-stationarity and correcting eventual forecasts bias while long windows will allow for a more accurate estimation of the errors distribution. To get ”the best of both worlds”, we proceed as in Lipiecki et al. (2024): We consider four different window sizes: 28 days (4 weeks), 56 days (8 weeks), 91 days (13 weeks) and 182 days (26 weeks) yielding four cdf estimates . We then compute an ensemble prediction considering vertical (probability) averaging (Lichtendahl et al., 2013):
@@ 2.5 Benchmark curves and price forecasting models
Ziel and Steinert (2016)
While our approach differs from that of Ziel and Steinert (2016) regarding the curves vector representation, the multivariate forecasting techniques applied to that representation are essentially the same. For that reason, to guarantee a fair comparison between the two approaches, we apply the models detailed in the previous section to (i) our vector of supply and demand FPCA scores (ii) the vector of supply and demand price class volumes obtained with the method described in Ziel and Steinert (2016). We invite the reader to refer to the authors’ paper for a rigorous and detailed description of this procedure (as well as for curves reconstruction), which we refer to as Ziel-Steinert Transformation (ZST), as opposed to our functional principal component analysis (FPCA).
@@
Price-based point forecasting models
We consider three benchmark price-based clearing price forecasting models: ARX, fARX and the Lasso-Estimated AutoRegressive model (LEAR) of Lago et al. (2021). The first two are already described in the previous section (equations 1 and 2). When both curve-based and price-based models are involved, to distinguish the clearing price forecasting model deriving from the curves model (curve-based) from that directly forecasting the price (price-based), we prefix the curve-based models with the curves representation used, e.g, FPCA-ARX or ZST-ARX.
Finally the LEAR model (equation 6) is very similar to fARX with the difference that it adds cross-hour dependence in the exogenous variables effects and accepts only two of them (usually the load and combined wind-solar generation day-ahead forecasts):
As in Lago et al. (2021), for all three models, prices are transformed with median-based scaling and the area hyperbolic sine variance stabilizing transformation.
Price-based probabilistic forecasting models
We test the three (price-based) clearing price probabilistic forecasting methods of Lipiecki et al. (2024), using the Julia package PostForecasts.jl (Lipiecki and Weron, 2025) developed by the authors: Quantile Regression Machine (QRM), Conformal Prediction (CP), Isotonic Distributional Regression (IDR). We also include the naive Normal (N) benchmark888Note that differently from Lipiecki et al. (2024) paper, we consider the version which does not suppose zero-mean errors and therefore accounts for potential forecasts bias.. As these methods are based on post-processing of point predictions, we use the best performing price-based model as a point predictor. For more information, we invite the reader to directly refer to Lipiecki et al. (2024).
@@ 2.6 Forecasts evaluation
As suggested in Ramsay and Silverman (2005), we measure curves forecasting accuracy with the squared correlation function:
@@
Additionally, to test whether a model significantly outperforms another, we run the Diebold-Mariano (DM) test (Diebold and Mariano, 1995) on the -norm of daily functional forecast errors considering (i) the average error across the 24 hours (ii) separate tests run on each of the 24 hours999Considering daily forecast errors avoids the problem of the strong intraday error autocorrelation, typical in day-ahead price forecasting, which would violate DM tests assumptions (Weron, 2014).. Two one-sided tests are run for each pair of models.
We evaluate clearing price point predictions with the mean absolute error (MAE), the root mean squared error (RMSE) and the relative MAE (rMAE) as defined in Lago et al. (2021).
@@
Similarly to the curves forecasts, we perform DM tests for both point and probabilistic forecasts considering average and per-hour daily errors. For point forecasting we consider the absolute () errors as in Lago et al. (2021) while we use the CRPS, as was done in Nowotarski and Weron (2018), for probabilistic forecasting.
@@ 2.7 Daily recalibration
As advised by Lago et al. (2021) and in order to simulate a live forecasting setting, the entire modeling pipeline is recalibrated on a daily basis using a rolling window of 364 days (52 weeks), since year appears to the best single-window choice for EPF according to Marcjasz et al. (2018) and Hubicka et al. (2019). This means that for every day of the testing period:
@@
For FPCA, we combine the elbow and threshold methods (see section 2.2) by taking the maximum of the two suggested number of FPCs to retain, considering a ratio of 99% for the threshold method. While automating the latter is straightforward, the elbow method relies on a qualitative visual assessment of the scree plot and requires therefore a more sophisticated approach. To automatically find this elbow point in the recalibration procedure, we use the popular knee-point detection method of Satopaa et al. (2011).
For ZST, Ziel and Steinert (2016) do not provide discussion on their choice for the number of price classes to consider. Since our goal is to compare the efficiency of the two curves representation strategies, we should ideally match it on a daily basis with the dimension picked by FPCA. For simplicity though, we take fixed and equal to their maximum value observed across the FPCA recalibration procedure, leaving therefore a slight advantage to ZST.